OpenStudy (anonymous):

Express 45 = x as a logarithmic equation. log45 = x log4x = 5 log54 = x log5x = 4 @TuringTest @AravindG @Compassionate @Loser66 @paki @aaronq

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i am gong to make a guess that \(45=x\) really means \[\huge 4^5=x\]

OpenStudy (aravindg):

a^x=y Then \[\log_a y=x\]

OpenStudy (misty1212):

besides the fact that \(4^5\) is an actual number, namely \(1024\) and so this really says \[1024=x\] if for so me possible reason you need to write it in a different forr \[b^y=x\ff \log_b(x)=y\]

OpenStudy (misty1212):

\[b^y=x\iff \log_b(x)=y\]

OpenStudy (anonymous):

so your @misty1212 answer was correct?

OpenStudy (anonymous):

so whats the answer @AravindG